40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 19 13:38:26 2021, gautam, Update, LSC, PRFPMI 
    Reply  Fri Feb 19 18:22:42 2021, rana, Update, LSC, PRFPMI sensiing matrix woes 
Message ID: 15822     Entry time: Fri Feb 19 13:38:26 2021     Reply to this: 15827
Author: gautam 
Type: Update 
Category: LSC 
Subject: PRFPMI 

I forgot that I had already done some investigation into recovering the PRFPMI lock after my work on the RF source. I don't really have any ideas on how to explain (or more importantly, resolve) the poor seperation of MICH and PRCL sensed in our 3f (but also 1f) photodiodes, see full thread here. Anyone have any ideas? I don't think my analysis (=code) of the sensing matrix can be blamed - in DTT, just looking the spectra of the _ERR_DQ channels for the various photodiodes while a ssingle frequency line is driving the PRM/BS suspension, there is no digital demod phase that decouples the MICH/PRCL peak in any of the REFL port photodiode spectra.

ELOG V3.1.3-